# GEM: A Gym for Agentic LLMs

### v1

@@ 4.1 Benchmarking RL algorithms for LLMs
Benchmarking has been critical for the progress of RL, with OpenAI-Gym providing standardized environments that enabled systematic evaluation of algorithms (Raffin et al., 2021; Huang et al., 2022). Following this paradigm, GEM offers a unified testbed for agentic LLMs, where prior work often relied on bespoke tasks that complicate fair comparison. We benchmark all baseline algorithms introduced in Section˜3.2 (GRPO, PPO [^5] PPO in this work generally refers to turn-level PPO instead of token-level PPO commonly seen in single-turn dialogue scenarios (Ouyang et al., 2022)., REINFORCE, ReBN) across eight GEM environments under a unified experimental protocol. All algorithms are implemented using Oat (Liu et al., 2024) with hyperparameters detailed in Appendix˜F. Results are evaluated by mean episode return, sample efficiency, and stability.
@@ 4.3 Tool-integration in math and question-answering tasks
Qwen3-4B
Base
(no tool)
Base + RL
(no tool,
single env)
Base + RL
(no tool,
mixed env)
Base + RL
(with tool,
single env)
Base + RL
(with tool,
mixed env)
@@ 5 Agent evaluation with GEM
In addition to RL training, GEM can serve as a unified evaluation interface to test LLM agents’ performance. In this section, we present two example use cases where we evaluate agents powered by strong LLMs (GPT-5 (OpenAI, 2025), Gemini-2.5-Pro (Gemini Team, 2025) and Claude-Sonnet-4 (Anthropic, 2025a)) on two complex tasks: database operation via model context protocol (MCP) (Anthropic, 2025b) and terminal interaction via docker containers, both of which have been added to GEM following Appendix˜A.
@@ D.3 Multi-agent environments
GEM supports multi-agent settings where multiple agents interact within the same environment. This capability enables the development of agents that can collaborate, compete, or simulate realistic interactions with other entities.
Interface design. GEM provides a MultiAgentEnv base class that extends the standard Gym API to support multiple agents. The step() and reset() functions operate on dictionaries keyed by agent identifiers:
@@
4observations, infos = env.reset() # Dict[agent_id -> observation]
@@
To implement a custom environment, users inherit from MultiAgentEnv and implement observe(agent) and _process_actions(actions). The framework handles agent lifecycle management and cumulative rewards tracking. Turn coordination is managed via AgentSelector, which supports two modes: sequential (agents act one at a time in round-robin order) and parallel (all agents act simultaneously). The selector determines which agents are active at each step and automatically advances turns, enabling flexible multi-agent interaction patterns without manual bookkeeping.

### v2

@@ 4.1 Benchmarking RL algorithms for LLMs
Benchmarking has been critical for the progress of RL, with OpenAI-Gym providing standardized environments that enabled systematic evaluation of algorithms (Raffin et al., 2021; Huang et al., 2022; 2024). Following this paradigm, GEM offers a unified testbed for agentic LLMs, where prior work often relied on bespoke tasks that complicate fair comparison. We benchmark all baseline algorithms introduced in Section˜3.2 (GRPO, PPO [^5] PPO in this work generally refers to turn-level PPO instead of token-level PPO commonly seen in single-turn dialogue scenarios (Ouyang et al., 2022)., REINFORCE, ReBN) across eight GEM environments under a unified experimental protocol. All algorithms are implemented using Oat (Liu et al., 2024) with hyperparameters detailed in Appendix˜F. Results are evaluated by mean episode return, sample efficiency, and stability.
@@ 4.3 Tool-integration in math and question-answering tasks
Qwen3-4B Base (no tool) Base + RL (no tool, single env) Base + RL (no tool, mixed env) Base + RL (with tool, single env) Base + RL (with tool, mixed env)
@@ 5 Agent evaluation with GEM
In addition to RL training, GEM can serve as a unified evaluation interface to test LLM agents’ performance. In this section, we present two example use cases where we evaluate agents powered by strong LLMs (GPT-5 (OpenAI, 2025), Gemini-2.5-Pro (Gemini Team, 2025) and Claude-Sonnet-4 (Anthropic, 2025b)) on two complex tasks: database operation via model context protocol (MCP) (Anthropic, 2025a) and terminal interaction via docker containers, both of which have been added to GEM following Appendix˜A.
@@ 6 Conclusions
Acknowledgment
This project was partially supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (Award Number: T1 251RES2514).
@@ D.3 Multi-agent environments
Many scenarios considered within reinforcement learning, such as adversarial games, social dilemmae and multi-party control systems require multiple agents by definition [Albrecht et al., 2024]. This introduces a whole host of different design decisions which a general multi-agent API should be able to embrace. By studying previous approaches to defining multi-agent APIs [Terry et al., 2021, Lanctot et al., 2019, Bou et al., 2023, Hu et al., 2023] and condensing them into a simple abstraction scheme, GEM also provides a high-level API for multi-agent LLM RL.
In doing so, GEM enables convenient development of agents that can collaborate, compete, or simulate realistic interactions with other agents/entities, adding comprehensiveness to its extensible platform.
Design. GEM’s multi-agent API is designed with the following principles in mind:
  • Minimal Prescriptions. The API only imposes constraints and methods when it is helpful to do so. This includes the AgentSelector class which allows for arbitrary definitions of turn order via overriding.
  • One Abstraction, One API. There should be only one way to realize the environment abstraction regardless of the environment’s properties.
API overview. GEM provides a MultiAgentEnv base class that extends the standard Gym API to support multiple agents. The step() and reset() functions operate on dictionaries keyed by agent identifiers:
@@
4observations, infos = env.reset() # Dict[agent_id, observation]
@@
To implement a custom environment, users inherit from MultiAgentEnv and implement observe(agent) and _process_actions(actions). The framework handles agent lifecycle management and cumulative rewards tracking. Turn coordination is managed via AgentSelector, which supports two modes: sequential (agents act one at a time in round-robin order) and parallel (all agents act simultaneously). The environment can then determine which agents are active at each step and automatically advance turns, enabling flexible multi-agent interaction patterns without manual bookkeeping.
@@ Appendix E Related works
We also note that there are early works [Abdulhai et al., 2023, Tajwar et al., 2025] incorporating text games as the evaluation toolkit or for language model fine-tuning. However, they did not focus on a standardized suite of RL training environments nor did they investigate different RL algorithms.